### v1

# A General Control-Theoretic Approach for Reinforcement Learning: Theory and Algorithms

Abstract
We devise a control-theoretic reinforcement learning approach to support direct learning of the optimal policy. We establish theoretical properties of our approach and derive an algorithm based on a specific instance of this approach. Our empirical results demonstrate the significant benefits of our approach.
Keywords: Reinforcement Learning, Control Theory, Policy Gradient Methods.
1 Introduction
For many years now, reinforcement learning (RL) has succeeded in solving a wide variety of decision-making problems and control for robotics [1, 2, 3, 4, 5]. Generally speaking, model-free methods [6, 7] often suffer from high sample complexity that can require an inordinate amount of samples, making them unsuitable for robotic applications where collecting large amounts of data is time-consuming, costly and potentially dangerous for the system and its surroundings [8, 9, 10, 11, 12]. On the other hand, model-based RL methods have been successful in demonstrating significantly reduced sample complexity and in outperforming model-free approaches for various decision making under uncertainty problems (see, e.g., [13, 14]). However, such model-based approaches can suffer from the difficulty of learning an appropriate model and from worse asymptotic performance than model-free approaches due to model bias from inherently assuming the learned system dynamics model accurately represents the true system environment (see, e.g., [15, 16, 17]).
In this paper we propose a novel form of RL that seeks to directly learn an optimal control policy for a general underlying (unknown) dynamical system and to directly apply the corresponding learned optimal control policy within the dynamical system. This general approach is in strong contrast to many traditional model-based RL methods that, after learning the system dynamics model which is often of high complexity and dimensionality, then use this system dynamics model to compute an approximate solution of a corresponding (stochastic) dynamic programming problem, often applying model predictive control (see, e.g., [18]). Our control-based RL (CBRL) approach instead directly learns the unknown parameters that derive, through control-theoretic means, an optimal control policy function from a family of control policy functions, often of much lower complexity and dimensionality, from which the optimal control policy is directly obtained.
The theoretical foundation and analysis of our CRBL approach is presented within the context of a general Markov decision process (MDP) framework that extends the family of policies associated with the classical Bellman operator to a family of control-policy functions mapping a vector of (unknown) parameters from a corresponding parameter set to a control policy which is optimal under those parameters, and that extends the domain of these control policies from a single state to span across all (or a large subset of) states, with the (unknown) parameter vector encoding global and local information that needs to be learned. Within the context of this MDP framework and our general CBRL approach, we establish theoretical results on convergence and optimality with respect to (w.r.t.) a CBRL contraction operator, analogous to the Bellman operator.
The basic idea of learning a form of a parameterized policy within an MDP framework to reduce sample complexity is not a new idea. One popular approach along these lines concerns policy gradient methods [19, 20], where gradient ascent of the value function in a space of policies is used, possibly together with projection to obtain an optimal policy [21]. These ideas have been further refined in neural network based policy optimization approaches (see, e.g., [21]) such as TRPO [22] and PPO [23]. It is important to note, however, that our CBRL approach is fundamentally different in several important ways. More specifically, we exploit control-theoretic methods to derive the optimal policy in terms of estimates of a few (global) parameters and directly learn these unknown parameters in an iterative manner based on observations from applying the optimal control policy of the current estimate of parameters. Meanwhile, within the context of our general CBRL approach, we establish a new control-policy-parameter gradient ascent theorem, analogous to the standard policy gradient theorem, together with a corresponding gradient ascent algorithm based on the (unknown) parameter vector.
With its foundation being optimal control, our CBRL approach is particularly suitable for dynamical systems and therefore the control of robotic systems. In addition to its established theoretical properties, numerical experiments of various simple robotics tasks demonstrate the effectiveness of our CBRL approach in reducing the number of samples needed, which is a key requirement for the application of learning and decision-making algorithms in real robotic systems.
We first present our general CBRL approach and then adapt this approach to a particular control-theoretic framework. Empirical results are presented next, followed by a discussion of limitations and future work. The appendix provides all proofs, experimental details, and additional results.
2 Control-Based Reinforcement Learning Approach
Consider a standard RL framework (see, e.g., [20, 24]) in which a decision-making agent interacts with a stochastic environment modeled as an MDP defined over a set of states ${{\mathbb{X}}}$, a set of actions ${\mathbb{U}}$, a transition probability kernel $\mathbb{P}$, and a reward function ${r}$ mapping state-action pairs to a bounded subset of ${\mathbb{R}}$. We denote a probability distribution on ${\mathbb{U}}$ as ${\mathbb{W}}$ (i.e., ${\mathbb{W}}$ is an element of $[0,1]^{{\mathbb{U}}}$ whose values sum to $1$). The agent seeks to determine a policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ comprising the sequence of control actions $u_{1},u_{2},\ldots$ that maximizes the corresponding discounted infinite-horizon stochastic dynamic programming (sDP) formulation expressed as
$\max_{u_{1},u_{2},\ldots}\mathbb{E}_{\mathbb{P}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}{r}(x_{t},u_{t})\bigg{]},$ (1)
where $x_{t}\in{{\mathbb{X}}}$ denotes the state of the system at time $t\in{\mathbb{Z}}_{+}$, $u_{t}\in{\mathbb{U}}$ the control action at time $t$, $\gamma\in(0,1)$ the discount factor, and $\mathbb{E}_{\mathbb{P}}:=\mathbb{E}_{x_{t+1}\sim\mathbb{P}(\cdot|x_{t},u)}$ defines expectation w.r.t. $\mathbb{P}(\cdot|x_{t},u)$. The stationary policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ defines a distribution of available control actions given the current state $x$, which reduces to a deterministic policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{U}}$ when the conditional distribution renders a constant action for state $x$; with abuse of notation, we write policy $\pi(x)$. Let ${{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$ denote the space of bounded real-valued functions over ${{\mathbb{X}}}\times{\mathbb{U}}$ with supremum norm, let $\Pi$ denote the set of all stationary policies, and let $\pi^{\ast}(x)$ denote the optimal stationary policy.
Our CBRL approach consists of exploiting control-theoretic methods to solve the general sDP formulation (1) in terms of the unknown system and control variables, and directly learning these unknown parameters over time through a corresponding iterative solution process. In particular, our general approach considers two key ideas: (i) extending the family of policies $\Pi$ associated with classical RL to a family of control-policy functions that map a parameter vector $p$ from a parameter set to a control policy that is optimal under $p$; and (ii) extending the domain of these control policies from a single state to span across all states in ${{\mathbb{X}}}$, with the parameter vector $p$ encoding global and local information (e.g., gravity and mass) that needs to be learned. More formally, let ${{\mathbb{S}}}$ denote the set of parameter vectors, and ${\mathbb{F}}$ the family of control-policy functions whose elements comprise functions ${\mathcal{G}}:p\mapsto{\mathcal{F}}_{p}$ that map a parameter vector $p\in{{\mathbb{S}}}$ to a control policy ${\mathcal{F}}_{p}:{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ that is optimal under $p$. For any $p\in{{\mathbb{S}}}$, the control-policy function ${\mathcal{G}}(\cdot)\in{\mathbb{F}}$ identifies a specific control policy ${\mathcal{F}}_{p}\in\Pi$ that provides the best rewards in expectation across all states $x\in{{\mathbb{X}}}$ from among all control-policy functions in ${\mathbb{F}}$. Such control policy mappings ${\mathcal{G}}:{{\mathbb{S}}}\rightarrow\Pi$ derive optimal control policies from vectors in the parameter set ${{\mathbb{S}}}$ through control-theoretic methods, with the goal of learning the unknown parameter vector $p\in{{\mathbb{S}}}$ through a corresponding iterative solution process. Hence, once the parameter vector is learned, the desired optimal control policy is realized.
In the remainder of this section, we first establish theoretical properties of our CBRL approach and then consider an algorithm to compute solutions based on our CBRL approach and its properties.
Convergence and Optimality. We consider an analog of the Bellman operator and related theoretical results on convergence and optimality. For each $Q$-function $q(x,u)\in{{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$, define a generic CBRL function $\tilde{q}:{{\mathbb{X}}}\times\Pi\rightarrow{\mathbb{R}}$ as $\tilde{q}(x,{\mathcal{F}}_{p}):=q(x,{\mathcal{F}}_{p}(x))$ where the control policy ${\mathcal{F}}_{p}$ is obtained from the control-policy function ${\mathcal{G}}$ derived from the parameter vector $p\in{{\mathbb{S}}}$, and thus evidently $\tilde{q}(x,{\mathcal{F}}_{p})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ with ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ denoting the space of bounded real-valued functions over ${{\mathbb{X}}}\times\Pi$ with supremum norm. We then define our CBRL operator ${{\mathbf{T}}}$ on ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ as
$({{\mathbf{T}}}\tilde{q})(x,{\mathcal{F}}_{p}):=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}.$ (2)
Our approach comprises a progression from the family ${\mathbb{F}}$ to the control-policy functions ${\mathcal{G}}\in{\mathbb{F}}$ to the control policies ${\mathcal{F}}_{p}\in\Pi$ to the control action ${\mathcal{F}}_{p}(x)\in{\mathbb{W}}$, upon applying ${\mathcal{F}}_{p}$ in state $x\in{{\mathbb{X}}}$. We next introduce an important assumption on the richness of the policy family ${\mathbb{F}}$ within this progression.
Assumption 2.1.
There exist a policy function ${\mathcal{G}}$ in the family ${\mathbb{F}}$ and a unique vector $p^{\prime}$ in the parameter set ${{\mathbb{S}}}$ such that, for any state $x\in{{\mathbb{X}}}$, $\pi^{\ast}(x)={\mathcal{F}}_{p^{\prime}}(x)={\mathcal{G}}(p^{\prime})(x)$.
Assumption 2.1 says that ${\mathbb{F}}$ is rich enough to include a global policy that coincides with the Bellman operator for each state. We then have the following formal result on the convergence of the operator ${{\mathbf{T}}}$ of our CBRL approach and the optimality of this convergence w.r.t. the Bellman equation [20, 24].
Lemma 2.1.
For any $\gamma\in(0,1)$, operator ${{\mathbf{T}}}$ in (2) is a contraction in the supremum norm. Suppose Assumption 2.1 holds for the family of policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$. Then, our CBRL approach with contraction operator ${{\mathbf{T}}}$ achieves the same optimal outcome as the Bellman operator.
From the first part of Lemma 2.1 and for any function $\tilde{q}\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$, the iterations ${{\mathbf{T}}}^{t}(\tilde{q})$ converge as $t\rightarrow\infty$ to $\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})$, the unique fixed point of the contraction operator ${{\mathbf{T}}}$, and $\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})$ satisfies
$\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}^{\prime}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}.$ (3)
The second part of Lemma 2.1 implies that, under the contraction operator ${{\mathbf{T}}}$ and Assumption 2.1, our CBRL approach is optimal by realizing the same unique fixed point of the Bellman operator. We note, however, that this optimality is achieved with great reduction in sample complexity due in part to another important difference with standard RL, namely the search of our CBRL approach is across all states $x\in{{\mathbb{X}}}$ to find a single optimal parameter vector $p$ that derives a single optimal control-policy function ${\mathcal{G}}\in{\mathbb{F}}$ which coincides with the Bellman equation for each state. Iterations w.r.t. the operator ${{\mathbf{T}}}$ of our CBRL approach then consist of improving the estimates of the parameter vector $p$ while applying the optimal control-policy function ${\mathcal{G}}$ derived from the current estimate of $p$.
Before addressing this iterative process within the context of policy gradient methods, we first relax the above conditions to consider families that do not satisfy Assumption 2.1 and policy functions that map a parameter vector to a control policy whose domain spans across a large subset of states (as opposed to all states) in ${{\mathbb{X}}}$ and renders asymptotically optimal (as opposed to globally optimal) rewards. Supposing ${\mathbb{F}}$ satisfies Assumption 2.1, consider a sequence of less rich families ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k}\subset{\mathbb{F}}$ of policy functions ${\mathcal{G}}^{(i)}\in{\mathbb{F}}_{i}$ derived from parameter vectors of the corresponding sets ${{\mathbb{S}}}_{i}$, and define the operators ${{\mathbf{T}}}_{i}:\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$ as in (2) for any function $\tilde{q}_{i}(x,{\mathcal{F}}_{p}^{(i)})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$, $i\in[k]$. Then, from Lemma 2.1, each operator ${{\mathbf{T}}}_{i}$ under parameter set ${{\mathbb{S}}}_{i}$ is a contraction in the supremum norm and converges to the unique fixed point $\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}_{p}^{(i)})$ that satisfies the corresponding version of (3), for all $x\in{{\mathbb{X}}}$ and $p\in{{\mathbb{S}}}_{i}$, $i\in[k]:=\{1,\ldots,k\}$. We then have the following formal result on the asymptotic optimality of our CBRL approach in such approximate settings.
Theorem 2.1.
Assume the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$. Consider a sequence of families of policy functions ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ such that $\bigcup_{i=1}^{k}{\mathbb{F}}_{i}\rightarrow{\mathbb{F}}$ as $k\rightarrow\infty$, with ${{\mathbb{S}}}$ and ${{\mathbb{S}}}_{i}$ respectively denoting the parameter sets corresponding to ${\mathbb{F}}$ and ${\mathbb{F}}_{i}$, $i\in[k]$. Then, $\sup_{x\in{{\mathbb{X}}}}\|\tilde{q}_{k}^{\prime}-\tilde{q}^{\prime}\|\rightarrow 0$ as $k\rightarrow\infty$.
One instance of a sequence of families ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ in Theorem 2.1 consists of piecewise-linear control policies of increasing richness (e.g., the class of canonical piecewise-linear functions [25]) w.r.t. finer and finer granularity of the control policy function space converging towards ${\mathbb{F}}$.
Control-Policy-Parameter Gradient Ascent. We now return to our CBRL iterative process in the context of policy gradient methods, building on our foregoing results to establish theoretical results analogous to the standard policy gradient theorem that establishes, for policies parameterized by $\theta$, a connection between the gradient of the value function w.r.t. $\theta$ and the gradient of the policy action w.r.t. $\theta$. This theoretical result does not apply to our CBRL approach because there is no direct connection between the parameter vector $p\in{{\mathbb{S}}}$ and the control policy action. Rather, under CBRL and for any vector $p\in{{\mathbb{S}}}$, the control-policy function ${\mathcal{G}}(p)\in{\mathbb{F}}$ identifies a particular control policy ${\mathcal{F}}_{p}\in\Pi$, and then the control policy applied in any state $x\in{{\mathbb{X}}}$ yields the control action taken.
Recall that the stationary control policy ${\mathcal{F}}_{p}:{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ defines a probability distribution over all available control actions $u\in{\mathbb{U}}$ given the current state $x\in{{\mathbb{X}}}$. Define ${\mathcal{F}}_{p,u}(x)$ to be the element of ${\mathcal{F}}_{p}(x)$ corresponding to action $u$ (i.e., ${\mathcal{F}}_{p,u}(x):=\mathbb{P}[u_{t}=u\,|\,x_{t}=x,{\mathcal{F}}_{p}]$), and correspondingly define $\tilde{q}_{u}(x,{\mathcal{F}}_{p}):=q_{u}(x,{\mathcal{F}}_{p}(x)):=q(x,u)$. Let $x_{0}$ denote the start state at time $t=0$ and $\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)$ the probability of going from state $x_{0}$ to state $x$ in $k$ steps under the control policy ${\mathcal{F}}_{p}$. Our general control-policy-parameter gradient ascent result is then formally expressed as follows.
Theorem 2.2.
Consider a family of control policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$ with contraction operator ${{\mathbf{T}}}$ in the form of (2). We then have
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})=\sum_{x\in{{\mathbb{X}}}}\>\sum_{k=0}^{\infty}\gamma^{k}\left(\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)\right)\>\sum_{u\in{\mathbb{U}}}\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}\tilde{q}_{u}(x,{{\mathcal{F}}}_{p}),$ (4)
where it is assumed that ${\mathcal{F}}_{p,u}(x)$ is differentiable w.r.t. ${\mathcal{F}}_{p}$ and ${\mathcal{F}}_{p}(x)$ is differentiable w.r.t. $p$, i.e., both $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ and $\partial{\mathcal{F}}_{p}(x)/\partial p$ exist.
The corresponding gradient ascent result for the special case of deterministic control policies follows directly from Theorem 2.2 with the conditional probability distribution ${\mathcal{F}}_{p,u}(x)$ given by the associated indicator function ${\mathbb{I}}_{{\mathcal{F}}_{p}(x)}(u)$, returning $1$ if and only if $u={\mathcal{F}}_{p}(x)$ and zero otherwise.
Following along similar lines to those of various forms of policy-gradient ascent methods based on the standard policy gradient theorem, we devise control-policy-parameter gradient ascent methods within the context of our CBRL approach based on Theorem 2.2. One such gradient ascent method comprises an iterative process for directly learning the unknown vector $p$ of the optimal control policy w.r.t. the value function $V$ whose iterations, with stepsize $\eta$, proceed according to
$p_{t+1}=p_{t}+\eta\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{t}),$ (5)
where $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is as given by (4). We note that the standard policy gradient method is a special case of (5) where the parameter vector $p$ is directly replaced by the policy $\pi$ (see, e.g., [21]).
3 Adaptation of Control-Based Reinforcement Learning Approach
[…]
$A^{\top}P+PA-PBR^{-1}B^{\top}P+\underline{Q}=0.$ (6)
The optimal control action $u_{t}^{\ast}$ at time $t$ is then obtained from the solution $P$ of the continuous-time ARE (CARE) as $u_{t}^{\ast}=-Kx_{t}$, where $K=R^{-1}B^{\top}P$.
Consider our CBRL approach within the context of the LQR control-theoretic framework where at time $t$ the system is in state $x_{t}$ and the parameter vector estimate is $p_{t}$. The optimal (deterministic) control policy is then given by ${\mathcal{F}}_{p}(x_{t})=-Kx_{t}$ via the corresponding solution of the CARE (6), together with the change of coordinates $\tilde{x}=x-x^{*}$ in general when the target $x^{*}$ is not the origin. We then update the parameter vector according to the iterative process (5) where $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is obtained from (4) of Theorem 2.2. In particular, the first partial derivative term on the right hand side of (4) is given by the standard policy gradient solution, and the second partial derivative term on the right hand side of (4) is obtained by differentiation of the CARE (6), which in turn renders [27]
$dP\tilde{A}+\tilde{A}^{\top}dP+(dZ+dZ^{\top}+d\underline{Q}+K^{\top}dRK)=0,$ (7)
where $\tilde{A}=A-BK$ and $dZ=P(dA-dBK)$.
Our CBRL approach under the LQR control-theoretic framework concerns the linear dynamics $\dot{x}=Ax+Bu$ where $A$ and $B$ contain elements of the unknown parameter vector $p$. By leveraging known basic information about the LQR control problem at hand, only a relatively small number of unknown parameters need to be learned. For robotics and physics applications where state variables are derivatives of each other w.r.t. time (e.g., position, velocity, acceleration, jerk), the corresponding rows in the $A$ matrix consist of a single $1$ and the corresponding rows in $B$ comprise zeros. We exploit this basic information to consider general matrix forms for $A$ and $B$ that reduce the number of unknown parameters to be learned. As a representative illustration, the system dynamics when there are 2 groups of such variables have the form given in (9) of Appendix B.
4 Experimental Results
In this section, we conduct numerical experiments to evaluate the performance of our CBRL approach of Section 2 adapted to the LQR control-theoretic framework according to Section 3, as summarized in Algorithm 1 of Appendix A.4. Our numerical experiments consider several classical control tasks from Gymnasium [28], including Cart Pole, Lunar Lander (Continuous), Mountain Car (Continuous), and Pendulum. We compare our CBRL method with three state-of-the-art RL algorithms, namely DQN [7] for discrete actions, DDPG [29] for continuous actions and PPO [23], together with a variant of PPO that solely replaces the nonlinear policy of PPO with a linear policy (since we know the optimal policy for some of the problems, e.g., Cart Pole, is linear). These baselines are selected as the state-of-the-art RL algorithms for solving the control tasks under consideration. Experimental details and additional results, including validation of the robustness of our CBRL approach, are provided in Appendix C both in general and for each control task.
Our CBRL approach depends upon the control-theoretic framework to which it is adapted, where we have chosen LQR as a representative example with the understanding that not all control tasks can be adequately solved using LQR even if the parameter vector $p$ is known. Recall, however, that our CBRL approach allows the domain of the control policies to span a subset of states in ${{\mathbb{X}}}$, thus enabling partitioning of the state space so that properly increased richness w.r.t. finer and finer granularity can provide improved approximations and asymptotic optimality according to Theorem 2.1, analogous to canonical piecewise-linear approximations [25]. While Cart Pole and Lunar Lander (Continuous) can be directly addressed within the context of LQR, this is not the case for Mountain Car (Continuous) and Pendulum which require a nonlinear controller. We therefore partition the state space in the case of such control tasks and consider a corresponding piecewise-LQR controller where the learned parameter vectors may differ or be shared across the partitions.
(a) Return v.s. Episode
(b) Return v.s. Time
(c) Parameter of CBRL
Figure 1: Learning curves of CartPole-v0 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DQN. (c): Learning behavior of CBRL parameters.
(a) Return v.s. Episode
(b) Return v.s. Time
(c) Parameter of CBRL
Figure 2: Learning curves of LunarLanderContinuous-v2 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
(a) Return v.s. Episode
(b) Return v.s. Time
(c) Parameter of CBRL
Figure 3: Learning curves of MountainCarContinuous-v0 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
(a) Return v.s. Episode
(b) Return v.s. Time
(c) Parameter of CBRL
Figure 4: Learning curves of Pendulum-v1 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
4.1 Cart Pole under CBRL LQR
Cart Pole consists of a pole connected to a horizontally moving cart with the goal of balancing the pole by applying a force on the cart to the left or right. The state of the system is $x=[s,\dot{s},\theta,\dot{\theta}]$ in terms of the position of the cart $s$, the velocity of the cart $\dot{s}$, the angle of the pole $\theta$, and the angular velocity of the pole $\dot{\theta}$. With upright initialization of the pole on the cart, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $195$.
The LQR controller can be used to solve the Cart Pole problem provided that all the parameters of the system are known (e.g., mass of cart, mass and length of pole, gravity). We address the problem within the context of our CBRL approach by exploiting the general matrix form for the LQR dynamics (10) (Appendix C.1), solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},\ldots,a_{7},b_{0},b_{1}$ to be learned (Fig. 1(c)).
Fig. 1 presents numerical results for the three state-of-the-art baselines (discrete actions) and our CBRL approach, with each run over five independent random seeds. All parameters in (10) are initialized uniformly in $(0,1)$ and then learned using our policy-parameter gradient ascent iteration (5). Fig. 1(a)–(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, as well as demonstrating a more stable training process.
4.2 Lunar Lander (Continuous) under CBRL LQR
Lunar Lander (Continuous) is a classical spacecraft trajectory optimization problem with the goal to land softly and fuel-efficiently on a landing pad by applying thrusters to the left, to the right and upward. The state of the system is $x=[s_{x},v_{x},s_{y},v_{y},\theta,\dot{\theta}]$ in terms of $(x,y)$ positions $s_{x}$ and $s_{y}$, two linear velocities $v_{x}$ and $v_{y}$, angle $\theta$, and angular velocity $\dot{\theta}$. Starting at the top center of the viewport with a random initial force applied to its center of mass, the lander (spacecraft) is subject to gravity, friction and turbulence while surfaces on the “moon” are randomly generated in each episode with the target landing pad centered at $(0,0)$. The problem is considered “solved” upon achieving an average return of $200$.
The LQR controller can be used to solve the Lunar Lander (Continuous) problem provided that all the parameters of the system are known (e.g., mass of lander, gravity, friction, etc). We address the problem within the context of our CBRL approach by exploiting the general matrix form for the LQR dynamics (11) (Appendix C.2), solely taking into account mild physical information from the system state, with the unknown parameters $a_{0},\ldots,a_{11},b_{0},b_{1},b_{2}$ to be learned (Fig. 2(c)).
Fig. 2 presents numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds. All parameters in (11) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). Fig. 2(a)–(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. We note that the baseline algorithms often crash and terminate sooner than the more successful landings of our CBRL approach, resulting in the significantly worse performance exhibited in Fig. 2(a)–(b).
4.3 Mountain Car (Continuous) under CBRL Piecewise-LQR
Mountain Car (Continuous) consists of the placement of a car in a valley with the goal of accelerating the car to reach the target at the top of the hill on the right by applying a force on the car to the left or right. The state of the system is $x=[s,v]$ in terms of the position of the car $s$ and the velocity of the car $v$. With random initialization at the bottom of the valley, the problem is considered “solved” upon achieving an average return of $90$.
Recalling that the LQR controller is not sufficient to solve this problem, we consider a piecewise-LQR controller that partitions the state space into two regions (see Fig. 9 in Appendix C.3). The target state $x^{*}=[s^{*},v^{*}]$ is selected to be $x^{*}=[-1.2,0]$ if $v<0$, and to be $x^{*}=[0.6,0]$ otherwise, where $-1.2$ and $0.6$ represent the position of the left hill and the right hill, respectively. We address the problem within the context of our CBRL approach by exploiting the general matrix form for the piecewise-LQR dynamics (12) (Appendix C.3), solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0}$ to be learned (Fig. 3(c)).
Fig. 3 presents numerical results for the three state-of-the-art (continuous actions) baselines and our CBRL approach, with each run over five independent random seeds. All parameters in (12) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). Fig. 3(a)–(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process.
4.4 Pendulum under CBRL Piecewise-LQR
Pendulum consists of a link attached at one end to a fixed point and the other end being free, with the goal of swinging up to an upright position by applying a torque on the free end. The state of the system is $x=[\theta,\dot{\theta}]$ in terms of the angle of the link $\theta$ and the angular velocity of the link $\dot{\theta}$. With random initial position of the link, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $-200$.
Recalling that the LQR controller is not sufficient to solve the Pendulum problem, we consider a piecewise-LQR controller that partitions the state space into four regions (see Fig. 11 in Appendix C.4). The target state $x^{*}=[\theta^{*},\dot{\theta}^{*}]$ is selected based on the boundary angle in each partition (counter-clockwise boundary angle if $\dot{\theta}>0$, and clockwise boundary angle otherwise), and the energy conservation law for target angular velocity (see Appendix C.4 for more details). We address the problem within the context of our CBRL approach by exploiting the general matrix form for the piecewise-LQR dynamics (13) (Appendix C.4), solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0},d_{0}$ to be learned (Fig. 4(c)).
Fig. 4 presents numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds. All parameters in (13) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). Fig. 4(a) illustrates a comparable performance between our CBRL approach and the best baseline DDPG in terms of the convergence w.r.t. the number of episodes, while Fig. 4(b) clearly demonstrates that CBRL outperforms all other baselines w.r.t. running time, in addition to demonstrating a more stable training process.
5 Discussion and Limitations
In this paper we devise a CBRL approach to directly learn the optimal policy, establishing its theoretical properties, adapting it to the LQR control-theoretic framework, and empirically demonstrating its performance benefits. We now turn to discuss limitations of our CBRL approach. Although primarily used for theoretical purposes, the supremum of our CBRL operator (2) and its fixed point (3) may be challenging if computed directly when the parameter space is continuous. This issue, however, is no different than corresponding challenges in standard RL when the action space is continuous, and thus similar steps can be taken to address the issue which does not arise in our Algorithm 1. The stepsize selection and parameter-vector initialization are also important implementation issues, just like in nonconvex optimization and RL. As described in Appendix C, the former issue can be addressed with similar steps taken in standard RL, including the adaptive stepsize selection in our implementation of Algorithm 1, and experimental results demonstrate the robustness of our CBRL approach w.r.t. the latter issue. Given the efficiency of our CBRL approach relative to state-of-the-art methods, additional computations for stepsize selection and parameter-vector initialization can be employed when issues are encountered while retaining overall computational efficiencies. Future work includes exploring other control-theoretic frameworks that can solve problems with different characteristics. For instance, consideration of Model Predictive Control as the control-theoretic framework instead of LQR would allow us to impose state constraints related to, e.g., safety.
Acknowledgments
This work was supported by IBM through the IBM-Rensselaer Future of Computing Research Collaboration.
Appendix A Control-Based Reinforcement Learning Approach
In this appendix, we present the proofs of our main theoretical results on control-based RL together with additional results and technical details related to Section 2. We also present the algorithmic details of our control-policy-parameter gradient ascent method used for the results in Section 4.
A.1 Proof of Lemma 2.1.
Lemma 2.1. For any $\gamma\in(0,1)$, operator ${{\mathbf{T}}}$ in (2) is a contraction in the supremum norm. Suppose Assumption 2.1 holds for the family of policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$. Then, our CBRL approach with contraction operator ${{\mathbf{T}}}$ achieves the same optimal outcome as the Bellman operator.
Proof.
For the operator ${{\mathbf{T}}}$ defined on ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ in (2) and for any two functions $\tilde{q}_{1}(x,{\mathcal{F}}_{p_{1}})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ and $\tilde{q}_{2}(x,{\mathcal{F}}_{p_{2}})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ with $p_{1},p_{2}\in{{\mathbb{S}}}$, we obtain
$\|{{\mathbf{T}}}\tilde{q}_{1}-{{\mathbf{T}}}\tilde{q}_{2}\|_{\infty}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\Big{|}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-r(x,{\mathcal{F}}_{p}(x))-\gamma\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{]}\Big{|}$
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\Big{|}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{]}\Big{|}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{|}\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{|}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\sup_{z\in{{\mathbb{X}}},p^{\prime}\in{{\mathbb{S}}}}\left|\tilde{q}_{1}(z,{\mathcal{F}}_{p^{\prime}})-\tilde{q}_{2}(z,{\mathcal{F}}_{p^{\prime}})\right|$
$\stackrel{{\scriptstyle(e)}}{{\leq}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\|\tilde{q}_{1}-\tilde{q}_{2}\|_{\infty}$
=$\stackrel{{\scriptstyle(f)}}{{=}}$ $\;\gamma\|\tilde{q}_{1}-\tilde{q}_{2}\|_{\infty},$
where (a) is by definition in (2), (b) follows from straightforward algebra, (c) and (d) are due to the triangle inequality, and (e) and (f) directly follow by definition. For any $\gamma\in(0,1)$, this establishes that the operator ${{\mathbf{T}}}$ in (2) is a contraction in the supremum norm, thus rendering the desired result for the first part.
For the second part of the lemma, under the stated supposition, we know that the optimal policy $\pi^{\ast}(x)$ realized by the Bellman operator holds for a unique vector $p^{\prime}$ in the parameter set ${{\mathbb{S}}}$ and any $x\in{{\mathbb{X}}}$. We also know that the Bellman operator and the operator ${{\mathbf{T}}}$ in (2) are both contractions in supremum norm with unique fixed points, where the fixed point of the Bellman operator is the optimal solution of the Bellman equation. By showing that these two fixed points coincide, the desired result follows.
To this end, from the definition of $\tilde{q}(x,{\mathcal{F}}_{p^{\prime}})$ and under the stated supposition, we have $\tilde{q}(x,{\mathcal{F}}_{p^{\prime}})=q(x,{\mathcal{F}}_{p^{\prime}}(x))=q(x,\pi^{\ast}(x))$ for any $x\in{{\mathbb{X}}}$, where $p^{\prime}$ is the unique vector in the parameter set ${{\mathbb{S}}}$ associated with $\pi^{\ast}(x)$. Then, from the right hand side of (2), we conclude $\tilde{q}(x,{\mathcal{F}}_{p^{\prime}})=q(x,{\mathcal{F}}_{p^{\prime}}(x))=q(x,\pi^{\ast}(x))\leq\sup_{p^{{}^{\prime\prime}}\in{{\mathbb{S}}}}\tilde{q}(x,{\mathcal{F}}_{p^{\prime\prime}})$. However, the last inequality must hold with equality due to the optimality of the policy $\pi^{\ast}(x)$, and therefore the fixed point of the operator ${{\mathbf{T}}}$ coincides with the fixed point of the Bellman operator, thus rendering the desired result. ∎
A.2 Proof of Theorem 2.1.
From Lemma 2.1, for $i\in[2]$, we have that the contraction operators ${{\mathbf{T}}}_{i}:{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi_{i})$ under the parameter sets ${{\mathbb{S}}}_{i}$ converge to the unique fixed points $\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}_{p}^{(i)})$ which satisfy, for all $x\in{{\mathbb{X}}}$ and $p\in{{\mathbb{S}}}_{i}$,
$\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}^{(i)}_{p})=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}^{(i)}_{p}(x)}(x,y)\bigg{[}r(x,{\mathcal{F}}^{(i)}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}_{i}}\tilde{q}_{i}^{\prime}(y,{\mathcal{F}}^{(i)}_{p^{\prime}})\bigg{]}.$ (8)
We first introduce the following lemma used in the proof of Theorem 2.1.
Lemma A.1.
Assume the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$. For the two parameter sets ${{\mathbb{S}}}_{1}$ and ${{\mathbb{S}}}_{2}$ above and any two parameter vectors $p_{1}\in{{\mathbb{S}}}_{1}$ and $p_{2}\in{{\mathbb{S}}}_{2}$, let $d(\cdot,\cdot)$ be a $\sup$-norm distance function defined over the action space ${\mathbb{U}}$, i.e., $d({\mathcal{F}}^{(1)}_{p_{1}},{\mathcal{F}}^{(2)}_{p_{2}})=\sup_{x\in{{\mathbb{X}}}}\|{\mathcal{F}}^{(1)}_{p_{1}}(x)-{\mathcal{F}}^{(2)}_{p_{2}}(x)\|$. Then, for all $\epsilon>0$ there exists $\delta>0$ such that, if for all $p_{1}\in{{\mathbb{S}}}_{1}$ there exists $p_{2}\in{{\mathbb{S}}}_{2}$ with $d({\mathcal{F}}^{(1)}_{p_{1}},{\mathcal{F}}^{(2)}_{p_{2}})<\delta$ and if for all $p_{2}\in{{\mathbb{S}}}_{2}$ there exists $p_{1}\in{{\mathbb{S}}}_{1}$ with $d({\mathcal{F}}^{(1)}_{p_{1}},{\mathcal{F}}^{(2)}_{p_{2}})<\delta$, we have $\sup_{x\in{{\mathbb{X}}}}\|\tilde{q}_{1}^{\prime}-\tilde{q}_{2}^{\prime}\|<\epsilon$.
Intuitively, Lemma A.1 shows that, for any policy families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$ sufficiently close to each other, the fixed points $\tilde{q}_{1}$ and $\tilde{q}_{2}$ of the corresponding operators ${{\mathbf{T}}}_{1}$ and ${{\mathbf{T}}}_{2}$ are also close to each other. When the policy families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$ are sufficiently rich and approach ${\mathbb{F}}$, then the fixed points of the corresponding operators ${{\mathbf{T}}}_{1}$ and ${{\mathbf{T}}}_{2}$ approach the unique fixed point of ${\mathbb{F}}$ satisfying (3), and therefore they approach the optimal $q$-value as promised by Bellman from Lemma 2.1. We formally characterize this asymptotic convergence of approximate optimality to global optimality in Theorem 2.1.
Theorem 2.1. Assume the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$. Consider a sequence of families of policy functions ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ such that $\bigcup_{i=1}^{k}{\mathbb{F}}_{i}\rightarrow{\mathbb{F}}$ as $k\rightarrow\infty$, with ${{\mathbb{S}}}$ and ${{\mathbb{S}}}_{i}$ respectively denoting the parameter sets corresponding to ${\mathbb{F}}$ and ${\mathbb{F}}_{i}$, $i\in[k]$. Then, $\sup_{x\in{{\mathbb{X}}}}\|\tilde{q}_{k}^{\prime}-\tilde{q}^{\prime}\|\rightarrow 0$ as $k\rightarrow\infty$.
Proof.
First, in the definition of the contraction mapping ${{\mathbf{T}}}$, the result of the supremum depends continuously on the set ${{\mathbb{S}}}$ and thus ${{\mathbf{T}}}$ depends continuously on ${{\mathbb{S}}}$. Next, since the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$, then [30, Lemma 2.2] shows that the unique fixed point depends continuously on the contraction mapping ${{\mathbf{T}}}$ and thus we find that $\tilde{q}^{\prime}$ depends continuously on the sets ${{\mathbb{S}}}_{i}$. This implies Lemma A.1, from which Theorem 2.1 then follows directly. ∎
A.3 Proof of Theorem 2.2.
Theorem 2.2. Consider a family of control policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$ with contraction operator ${{\mathbf{T}}}$ in the form of (2). We then have
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})=\sum_{x\in{{\mathbb{X}}}}\>\sum_{k=0}^{\infty}\gamma^{k}\left(\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)\right)\>\sum_{u\in{\mathbb{U}}}\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}\tilde{q}_{u}(x,{{\mathcal{F}}}_{p}),$ (4)
where it is assumed that ${\mathcal{F}}_{p,u}(x)$ is differentiable w.r.t. ${\mathcal{F}}_{p}$ and ${\mathcal{F}}_{p}(x)$ is differentiable w.r.t. $p$, i.e., both $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ and $\partial{\mathcal{F}}_{p}(x)/\partial p$ exist.
Proof.
We derive
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})$
=$\stackrel{{\scriptstyle(a)}}{{=}}\nabla_{p}\left[\sum_{u}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{u}\left[\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+{\mathcal{F}}_{p,u}(x_{0})\nabla_{p}\tilde{q}_{u}(x,{\mathcal{F}}_{p})\right]$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{u}\Bigg{[}\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\nabla_{p}\sum_{x^{\prime},r}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime},r|x,u)(r+V_{{\mathcal{F}}_{p}}(x^{\prime}))\Bigg{]}$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{u}\left[\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\sum_{x^{\prime}}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime}|x,u)\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime})\right]$
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{u}\Bigg{[}\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\sum_{x^{\prime}}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime}|x,u)\sum_{u^{\prime}}\bigg{[}\nabla_{p}{\mathcal{F}}_{p,u^{\prime}}(x^{\prime})\tilde{q}_{u}(x^{\prime},{\mathcal{F}}_{p})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\gamma{\mathcal{F}}_{p,u^{\prime}}(x^{\prime})\sum_{x^{\prime\prime}}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime\prime}|x^{\prime},u^{\prime})\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime\prime})\bigg{]}\Bigg{]}$
=$\stackrel{{\scriptstyle(f)}}{{=}}\sum_{x\in{{\mathbb{X}}}}\sum_{k=0}^{\infty}\gamma^{k}\left(\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)\right)\sum_{u}\nabla_{p}{\mathcal{F}}_{p,u}(x)\tilde{q}_{u}(x,{\mathcal{F}}_{p}),$
where (a) is by the definition of the value function for state $x_{0}$, (b) follows from the product rule, (c) follows by the definition of $\tilde{q}_{u}(x,{\mathcal{F}}_{p})$, (d) follows by applying the gradient w.r.t. $p$ summed over all $r$, (e) follows by repeating each of the preceding steps for $\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime})$, and (f) follows from repeated unrolling along the lines of (e) and upon recalling $\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)$ to be the probability of going from state $x_{0}$ to state $x$ in $k$ steps under the control policy ${\mathcal{F}}_{p}$.
The result then follows since $\nabla_{p}{\mathcal{F}}_{p,u}(x)=\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}$, by the chain rule. ∎
A.4 Control-Policy-Parameter Gradient Ascent Algorithm
Based on our new CBRL gradient ascent theorem (Theorem 2.2), we devise control-policy-parameter gradient ascent methods within the context of our general CBRL approach. One such gradient ascent method for directly learning the unknown parameter vector $p$ of the optimal control policy w.r.t. the value function $V$ comprises the iterative process according to (5) with stepsize $\eta$. Here $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is as given by (4), where the first gradient term $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ is essentially the standard policy gradient, and the second gradient term $\partial{\mathcal{F}}_{p}(x)/\partial p$ is specific to the control-theoretic framework employed in our general CBRL approach; refer to Section 3. We note that the standard policy gradient ascent method is a special case of (5) where the parameter vector $p$ is directly replaced by the policy $\pi$; see, e.g., [21].
Our algorithmic implementation of the above control-policy-parameter gradient ascent method is summarized in Algorithm 1. This algorithm, together with the adaptation of our general CBRL approach to the LQR control-theoretic framework according to Section 3, is used to obtain the numerical results for our CBRL approach in Section 4.
Algorithm 1 Control-Policy-Parameter Gradient Ascent Algorithm
$\mathcal{E},\mathbb{X}$, $\mathbb{U},\mathbb{F},{\mathcal{F}}_{p},\mathbf{u}(\cdot,\cdot),D_{0}=\Phi,\eta,N$triangleright Environment, State Set, Control Action Set, Control-Policy Family, Control-Policy, Control Action Generator (CBRL-LQR), Empty Rollout Dataset, Stepsize, Number of Iterations
Initialize $p_{0}\mid{\mathcal{F}}_{p_{0}}\in\mathbb{F}$, $\mathbf{X}_{0}\in\mathbb{X}$
for $t=0,1,\cdots,N$ do
    1. Generate the control action given the current state and control-policy-parameter
$\mathbf{u_{t}}=\mathbf{u}(\mathbf{X}_{t},p_{t})$
    2. Collect the rollout transition data by applying the control action to the environment $\mathcal{E}$
$\mathcal{E}\stackrel{{\scriptstyle\mathbf{u_{t}}}}{{\longrightarrow}}\textbf{rollout}:=(\mathbf{X}_{t},\mathbf{u_{t}},{r},\mathbf{X}_{t+1})$
$D_{t+1}=D_{t}\cup\textbf{rollout}$
    3. Compute the control-policy-parameter gradient (4) given the current rollout dataset $D_{t+1}$
    4. Update the control-policy-parameter using the gradient ascent iteration (5)
$p_{t+1}=p_{t}+\eta\nabla_{p}V_{{{\mathcal{F}}_{p}}}(\mathbf{X}_{t})$
end for
Appendix B Adaptation of Control-Based Reinforcement Learning Approach
Our CBRL approach under the LQR control-theoretic framework concerns the linear dynamics $\dot{x}=Ax+Bu$ where $A$ and $B$ contain elements of the unknown parameter vector $p$. By leveraging known basic information about the LQR control problem at hand, only a relatively small number of unknown parameters need to be learned. For robotics and physics applications where state variables are derivatives of each other w.r.t. time (e.g., position, velocity, acceleration, jerk), the corresponding rows in the $A$ matrix consists of a single $1$ and the corresponding rows in $B$ comprise zeros. We exploit this basic information to consider general matrix forms for $A$ and $B$ that reduce the number of unknown parameters to be learned. As a representative illustration, the system dynamics when there are two groups of such variables have the form given by
$\underbrace{\begin{bmatrix}\frac{dx_{1}}{dt}\\ \vdots\\ \frac{d^{k_{1}}x_{1}}{dt^{k_{1}}}\\ \frac{dx_{2}}{dt}\\ \vdots\\ \frac{d^{k_{2}}x_{2}}{dt^{k_{2}}}\end{bmatrix}}_{\dot{x}}=\underbrace{\begin{bmatrix}0&1&0&\cdots&&&0\\ 0&0&1&\cdots&&\\ a_{10}&\cdots&&&&&a_{1n}\\ 0&\cdots&&0&1&\dots&0\\ 0&\cdots&&&0&1&\cdots\\ a_{20}&\cdots&&&&&a_{2n}\end{bmatrix}}_{A}x+\underbrace{\begin{bmatrix}0\\ \vdots\\ b_{0}\\ 0\\ \vdots\\ b_{1}\end{bmatrix}}_{B}u.$ (9)
Appendix C Experimental Results
In this appendix we provide additional details and results based on our numerical experiments to evaluate the performance of our CBRL approach of Section 2 adapted to the LQR control-theoretic framework according to Section 3 and using Algorithm 1. Recall that we consider the following classical control tasks from Gymnasium [28]: Cart Pole, Lunar Lander (Continuous), Mountain Car (Continuous), and Pendulum. Our CBRL approach is compared against the three state-of-the-art RL algorithms DQN [7] for discrete actions, DDPG [29] for continuous actions and PPO [23], together with a variant of PPO that solely replaces the nonlinear policy of PPO with a linear policy (since we know the optimal policy for problems such as Cart Pole is linear). These baselines are selected as the state-of-the-art RL algorithms for solving the control tasks under consideration.
Our CBRL approach depends upon the control-theoretic framework to which it is adapted, where we have chosen LQR as a representative example with the understanding that not all control tasks can be adequately solved using LQR even if the parameter vector $p$ is known. Recall, however, that our CBRL approach allows the domain of the control policies to span a subset of states in ${{\mathbb{X}}}$, thus enabling the partitioning of the state space so that properly increased richness w.r.t. finer and finer granularity can provide improved approximations and asymptotic optimality according to Theorem 2.1 of Section 2, analogous to the class of canonical piecewise-linear approximations [25]. While Cart Pole and Lunar Lander (Continuous) can be directly addressed within the context of LQR, this is not the case for Mountain Car (Continuous) and Pendulum which require a nonlinear controller. We therefore partition the state space in the case of these control tasks and consider a corresponding piecewise-LQR controller where the learned parameter vectors may differ or be shared across the partitions. Such details are provided below for Mountain Car (Continuous) and Pendulum.
C.1 Cart Pole under CBRL LQR
Environment.
As depicted in Figure 5, Cart Pole consists of a pole connected to a horizontally moving cart with the goal of balancing the pole by applying a force on the cart to the left or right. The state of the system is $x=[s,\dot{s},\theta,\dot{\theta}]$ in terms of the position of the cart $s$, the velocity of the cart $\dot{s}$, the angle of the pole $\theta$, and the angular velocity of the pole $\dot{\theta}$. With upright initialization of the pole on the cart, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $195$.
Figure 5: The CartPole-v0 environment.
Controller.
The LQR controller can be used to solve the Cart Pole problem provided that all the parameters of the system are known (e.g., mass of cart, mass and length of pole, gravity), where the angle is kept small by our CBRL approach adapted to the LQR framework. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the LQR dynamics
$\dot{x}=\begin{bmatrix}\dot{s}\\ \ddot{s}\\ \dot{\theta}\\ \ddot{\theta}\end{bmatrix}=\underbrace{\begin{bmatrix}0&1&0&0\\ a_{0}&a_{1}&a_{2}&a_{3}\\ 0&0&0&1\\ a_{4}&a_{5}&a_{6}&a_{7}\end{bmatrix}}_{A}x+\underbrace{\begin{bmatrix}0\\ b_{0}\\ 0\\ b_{1}\end{bmatrix}}_{B}u,$ (10)
which solely takes into account general physical relationships (e.g., the derivative of the angle of the pole is equivalent to its angular velocity) and laws (e.g., the force can only affect the acceleration) with the unknown parameters $a_{0},\ldots,a_{7},b_{0},b_{1}$ to be learned.
Numerical Results.
Figure 6 presents our numerical results for the three state-of-the-art baselines (discrete actions) and our CBRL approach, each run over five independent random seeds. All parameters in (10) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 1) to validate the robustness of our CBRL approach. Figure 6(a) and 6(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. Figure 6(c) – 6(f) shows the learning behavior of CBRL parameters given different initialization.
Table 1: Initial Parameters of CartPole-v0
Name $a_{0}$ $a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ $a_{5}$ $a_{6}$ $a_{7}$ $b_{0}$ $b_{1}$
Initial Parameters 1 (P1) 0.436 0.026 0.55 0.435 0.42 0.33 0.205 0.619 0.3 0.267
Initial Parameters 2 (P2) 0.076 0.78 0.438 0.723 0.978 0.538 0.501 0.072 0.268 0.5
Initial Parameters 3 (P3) 0.154 0.74 0.263 0.534 0.015 0.919 0.901 0.033 0.957 0.137
Initial Parameters 4 (P4) 0.295 0.531 0.192 0.068 0.787 0.656 0.638 0.576 0.039 0.358
(a) Return v.s. Episode
(b) Return v.s. Time
(c) P1
(d) P2
(e) P3
(f) P4
Figure 6: Learning curves of CartPole-v0 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DQN. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 1.
C.2 Lunar Lander (Continuous) under CBRL LQR
Environment.
As depicted in Figure 7, Lunar Lander (Continuous) is a classical spacecraft trajectory optimization problem with the goal to land softly and fuel-efficiently on a landing pad by applying thrusters to the left, to the right, and upward. The state of the system is $x=[s_{x},v_{x},s_{y},v_{y},\theta,\dot{\theta}]$ in terms of $(x,y)$ positions $s_{x}$ and $s_{y}$, two linear velocities $v_{x}$ and $v_{y}$, angle $\theta$, and angular velocity $\dot{\theta}$. Starting at the top center of the viewport with a random initial force applied to its center of mass, the lander (spacecraft) is subject to gravity, friction and turbulence while surfaces on the “moon” are randomly generated in each episode with the goal landing pad centered at $(0,0)$. The problem is considered “solved” upon achieving an average return of $200$.
Figure 7: The LunarLanderContinuous-v2 environment.
Controller.
The LQR controller can be used to solve the Lunar Lander (Continuous) problem provided that all parameters of the system are known (e.g., mass of lander, gravity, friction, etc), where the angle is kept small by our CBRL approach adapted to the LQR framework. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the LQR dynamics
$\dot{x}=\begin{bmatrix}v_{x}\\ \ddot{s}_{x}\\ v_{y}\\ \ddot{s}_{y}\\ \dot{\theta}\\ \ddot{\theta}\end{bmatrix}=\underbrace{\begin{bmatrix}0&1&0&0&0&0\\ 0&a_{0}&0&a_{1}&a_{2}&a_{3}\\ 0&0&0&1&0&0\\ 0&a_{4}&0&a_{5}&a_{6}&a_{7}\\ 0&0&0&0&0&1\\ 0&a_{8}&0&a_{9}&a_{10}&a_{11}\\ \end{bmatrix}}_{A}x+\underbrace{\begin{bmatrix}0&0\\ 0&b_{0}\\ 0&0\\ b_{1}&0\\ 0&0\\ 0&b_{2}\\ \end{bmatrix}}_{B}u,$ (11)
which solely takes into account general physical relationships (akin to Cart Pole) and mild physical information from the system state (e.g., the acceleration is independent of the position) with the unknown parameters $a_{0},\ldots,a_{11},b_{0},b_{1},b_{2}$ to be learned.
Numerical Results.
Figure 8 presents our numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (11) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 2) to validate the robustness of our CBRL approach. Figure 8(a) and 8(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time. Figure 8(c) – 8(f) shows the learning behavior of CBRL parameters given different initialization.
Table 2: Initial Parameters of LunarLanderContinuous-v2
Name $a_{0}$ $a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ $a_{5}$ $a_{6}$ $a_{7}$
Initial Parameters 1 (P1) 0.551 0.708 0.291 0.511 0.893 0.896 0.126 0.207
Initial Parameters 2 (P2) 0.873 0.969 0.869 0.531 0.233 0.011 0.43 0.402
Initial Parameters 3 (P3) 0.778 0.238 0.824 0.966 0.973 0.453 0.609 0.776
Initial Parameters 4 (P4) 0.65 0.505 0.879 0.182 0.852 0.75 0.666 0.988
Name $a_{8}$ $a_{9}$ $a_{10}$ $a_{11}$ $b_{0}$ $b_{1}$ $b_{2}$
Initial Parameters 1 (P1) 0.051 0.441 0.03 0.457 0.649 0.278 0.676
Initial Parameters 2 (P2) 0.523 0.478 0.555 0.543 0.761 0.712 0.62
Initial Parameters 3 (P3) 0.642 0.722 0.035 0.298 0.059 0.857 0.373
Initial Parameters 4 (P4) 0.257 0.028 0.636 0.847 0.736 0.021 0.112
(a) Return v.s. Episode
(b) Return v.s. Time
(c) P1
(d) P2
(e) P3
(f) P4
Figure 8: Learning curves of LunarLanderContinuous-v2 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 2.
C.3 Mountain Car (Continuous) under CBRL Piecewise-LQR
Environment.
As depicted in Figure 9, Mountain Car (Continuous) consists of a car placed in a valley with the goal of accelerating the car to reach the goal on top of the right hill by applying a force on the car to the left or right. The state of the system is $x=[s,v]$ in terms of the position of the car $s$ and the velocity of the car $v$. With random initialization at the bottom of the valley, the problem is considered “solved” upon achieving an average return of $90$.
Figure 9: The MountainCarContinuous-v0 environment and partitions for piecewise-LQR.
Controller.
Recall that the LQR controller is not sufficient to solve the Mountain Car (Continuous) problem, even if all the parameters of the system are known (e.g., mass of the car and gravity), because a nonlinear controller is required. Consequently, we consider a piecewise-LQR controller that partitions the state space into two regions: LQR 1 and LQR 2 (see Figure 9). The target state $x^{*}=[s^{*},v^{*}]$ for LQR 1 and LQR 2 are respectively selected as $x^{*}=[-1.2,0]$ and $x^{*}=[0.6,0]$, where $-1.2$ and $0.6$ represent the position of the left hill and the right hill. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the piecewise-LQR dynamics
$\dot{e}=\begin{bmatrix}v-v^{*}\\ \dot{v}-\dot{v}^{*}\end{bmatrix}=\begin{bmatrix}0&1\\ a_{0}\sin(3s^{*})&a_{1}\end{bmatrix}e+\begin{bmatrix}0\\ b_{0}\end{bmatrix}(u-c_{0}\cos(3s^{*})),$ (12)
which solely takes into account general physical relationships and laws with the unknown parameters $a_{0},a_{1},b_{0},c_{0}$ to be learned, where $e=x-x^{*}$ and we select $v^{*}=0,\dot{v}^{*}=0$, and $s^{*}=-1.2$ or $0.6$.
Numerical Results.
Figure 10 presents our numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (12) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 3) to validate the robustness of our CBRL approach. Figure 10(a) and 10(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. Figure 10(c) – 10(f) shows the learning behavior of CBRL parameters given different initialization.
Table 3: Initial Parameters of MountainCarContinuous-v0
Name $a_{0}$ $a_{1}$ $b_{0}$ $c_{0}$
Initial Parameters 1 (P1) 0.549 0.715 0.603 0.545
Initial Parameters 2 (P2) 0.222 0.871 0.207 0.919
Initial Parameters 3 (P3) 0.771 0.021 0.634 0.749
Initial Parameters 4 (P4) 0.588 0.898 0.892 0.816
(a) Return v.s. Episode
(b) Return v.s. Time
(c) P1
(d) P2
(e) P3
(f) P4
Figure 10: Learning curves of MountainCarContinuous-v0 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 3.
C.4 Pendulum under CBRL Piecewise-LQR
Environment.
As depicted in Figure 11, Pendulum consists of a link attached at one end to a fixed point and the other end being free, with the goal of swinging up to an upright position by applying a torque on the free end. The state of the system is $x=[\theta,\dot{\theta}]$ in terms of the angle of the link $\theta$ and the angular velocity of the link $\dot{\theta}$. With random initial position of the link, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $-200$.
Figure 11: The Pendulum-v1 environment and partitions for piecewise-LQR.
Controller.
Recall that the LQR controller is not sufficient to solve the Pendulum problem, even if all the parameters of the system are known (e.g., mass of the link $m$, length of the link $l$, moment of inertia of the link $J$, and gravity $g$), because a nonlinear controller is required. Consequently, we consider a piecewise-LQR controller that partitions the state space into four regions: LQR 1, LQR 2, LQR 3, LQR 4 (see Figure 11). In terms of the target state $x^{*}=[\theta^{*},\dot{\theta}^{*}]$, $\theta^{*}$ is selected based on the boundary angle in each partition (counter-clockwise boundary angle if $\dot{\theta}>0$, and clockwise boundary angle otherwise), while $\dot{\theta}^{*}$ is selected based on the energy conservation law. More specifically, $\theta^{*}$ in LQR 1 is selected to be $\theta^{*}=\pi/4$ if $\dot{\theta}>0$, and to be $\theta^{*}=-\pi/4$ otherwise. Given the assumption that the link reaches a zero velocity at the upright position, $\dot{\theta}^{*}$ is selected at any position to be $\dot{\theta}^{*}=\sqrt{mgl(1-\cos(\theta))/J}$. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the piecewise-LQR dynamics
$\dot{e}=\begin{bmatrix}\dot{\theta}-\dot{\theta}^{*}\\ \ddot{\theta}-\ddot{\theta}^{*}\end{bmatrix}=\begin{bmatrix}0&1\\ a_{0}\cos(\theta^{*})&a_{1}\end{bmatrix}e+\begin{bmatrix}0\\ b_{0}\end{bmatrix}(u+c_{0}\sin(\theta^{*})),\quad\dot{\theta}^{*}=\sqrt{d_{0}(1-\cos(\theta))},$ (13)
which solely takes into account general physical relationships and laws with the unknown parameters $a_{0},a_{1},b_{0},c_{0},d_{0}$ to be learned, where $e=x-x^{*}$ and we select $\ddot{\theta}^{*}=0$.
Numerical Results.
Figure 12 presents our numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (13) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 4) to validate the robustness of our CBRL approach. Figure 12(a) illustrates a comparable performance between our CBRL approach and the best baseline DDPG in terms of the convergence w.r.t. the number of episodes, while Figure 12(b) clearly demonstrates that CBRL outperforms all other baselines w.r.t. running time, in addition to demonstrating a more stable training process. Figure 12(c) – 12(f) shows the learning behavior of CBRL parameters given different initialization.
Table 4: Initial Parameters of Pendulum-v1
Name $a_{0}$ $a_{1}$ $b_{0}$ $c_{0}$ $d_{0}$
Initial Parameters 1 (P1) 0.417 0.72 0.302 0.147 0.092
Initial Parameters 2 (P2) 0.893 0.332 0.821 0.042 0.108
Initial Parameters 3 (P3) 0.18 0.019 0.463 0.725 0.42
Initial Parameters 4 (P4) 0.223 0.523 0.551 0.046 0.361
(a) Return v.s. Episode
(b) Return v.s. Time
(c) P1
(d) P2
(e) P3
(f) P4
Figure 12: Learning curves of Pendulum-v1 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 4.
C.5 Implementation
Baselines.
As previously noted, the evaluation of our CBRL approach includes comparisons against the three baselines of DQN for discrete actions, DDPG for continuous actions and PPO, which are selected as the state-of-the-art RL algorithms for solving the control tasks under consideration. The additional baseline of replacing the nonlinear policy of PPO with a linear policy was again included because the optimal policy for some problems, e.g., Cart Pole, is known to be linear. This Linear algorithm solely replaces the policy network of PPO with a linear policy while maintaining all other components unchanged. More specifically, we decrease the number of hidden layers in the policy network by $1$, reduce the hidden layer size of the policy network to $16$, and remove all (nonlinear) activation functions in the policy network.
Stepsize.
The stepsize selection is an important implementation issue, just like in nonconvex optimization and RL. We address this issue by using similar methods employed in standard RL, i.e., the adaptive stepsize selection in our implementation of Algorithm 1. More concretely, we simply reduce the stepsize $\eta$ in (5) by a factor of $0.99$ each time achieving the “solved” return.
Hyperparameters
We summarize in Table 5 the hyperparameters that have been used in our experiments. To ensure a fair comparison, we maintain consistent hyperparameters (e.g., hidden layer size of the value network and the activation function) across all algorithms wherever applicable.
Table 5: Summary of Experimental Hyperparameters
Algorithm Linear PPO DDPG DQN CBRL
Action Generator NN NN NN NN CBRL-LQR
# of Hidden Layers (Policy) 1 2 2 2 N/A
# of Hidden Layers (Value) 2 2 2 2 2
Hidden Layer Size (Policy) 16 128 128 128 N/A
Hidden Layer Size (Value) 128 128 128 128 128
Activation (Policy) N/A ReLU & Tanh ReLU & Tanh ReLU & Tanh N/A
Activation (Value) ReLU ReLU ReLU ReLU ReLU
Discount Factor 0.99 0.99 0.99 0.99 0.99
Clip Ratio 0.2 0.2 N/A N/A N/A
Soft Update Ratio N/A N/A 0.005 N/A N/A